KyledaGreat:

Among all rectangles that have a perimeter of 64, find the dimensions of the one whose area is largest. Write your answers as fractions reduced to lowest terms. ____________, _____________

KyledaGreat:

@vocaloid

darkknight:

What math class is this?

KyledaGreat:

Algebra, this is Quadratic Functions

KyledaGreat:

Intermediate Algebra

darkknight:

Okay, just making sure it isn't calculus. So since this is not calculus you do guess and check to find the dimensions with the largest area. You will find a square to have a larger area than say, a rectangle of the same perimeter. It would be pretty cool if you try this for yourself as well. Plug in 16 and 16 for length and width, 20 and 12 for length and width (respectively), a few more length and widths that have a total Perimeter of 64, and see which has the greatest Area. Remember: A = L* W P = 2(L+W)

KyledaGreat:

P = 2(16+12)

darkknight:

?

darkknight:

Plug these in for L and W respectively (16, 16) (12, 20) (2, 30) (15, 17) into the area formula, I have already checked and the perimeter for each of these is 64. See which of these had the greatest Area

KyledaGreat:

like this: A = L* W A = 16 * 16 ?

darkknight:

yes, now 16 times 16 is....?

KyledaGreat:

256

darkknight:

yes, now do the same for the other coordinates

KyledaGreat:

12 * 20 = 240 2 * 30 = 60 15 *17 = 255

darkknight:

good job, so you found that 16*16 gives the largest area

darkknight:

So your answer, or the largest rectangles dimensions, are (16, 16) write 16 over both or the underlines in the question

KyledaGreat:

i could only do it like this

darkknight:

?

KyledaGreat:

1 attachment
darkknight:

Yes, you can submit

KyledaGreat:

ok

KyledaGreat:

could you answer more for me

darkknight:

Make a new post,